# An exploratory study on automatic identification of assumptions in the development of deep learning frameworks

### 08 Jan 2024 v1

@@ Abstract
Stakeholders constantly make assumptions in the development of deep learning (DL) frameworks. These assumptions are related to various types of software artifacts (e.g., requirements, design decisions, and technical debt) and can turn out to be invalid, leading to system failures. Existing approaches and tools for assumption management usually depend on manual identification of assumptions. However, assumptions are scattered in various sources (e.g., code comments, commits, pull requests, and issues) of DL framework development, and manually identifying assumptions has high costs (e.g., time and resources). To overcome the issues of manually identifying assumptions in DL framework development, we constructed a new and largest dataset (i.e., AssuEval) of assumptions collected from the TensorFlow and Keras repositories on GitHub; explored the performance of seven traditional machine learning models (e.g., Support Vector Machine, Classification and Regression Trees), a popular DL model (i.e., ALBERT), and a large language model (i.e., ChatGPT) of identifying assumptions on the AssuEval dataset. The experiment results show that: ALBERT achieves the best performance (f1-score: 0.9584) of identifying assumptions on the AssuEval dataset, which is much better than the other models (the 2nd best f1-score is 0.6211, achieved by ChatGPT). Though ChatGPT is the most popular large language model, we do not recommend using it to identify assumptions in DL framework development because of its low performance on the task. Fine-tuning ChatGPT specifically for assumption identification could improve the performance. This study provides researchers with the largest dataset of assumptions for further research (e.g., assumption classification, evaluation, and reasoning) and helps practitioners better understand assumptions and how to manage them in their projects.
@@ 1 Introduction
The importance of assumptions and their management in software development (including DL framework development) has been highlighted in many studies. For example, Corbató Corbato1991 mentioned in his ACM Turing Award lecture that “design bugs are often subtle and occur by evolution with early assumptions being forgotten as new features or uses are added to systems.” Garlan et al. pointed out that incompatible assumptions in software architecture can cause architectural mismatch Garlan2009 . Lewis et al. also got similar results in ML systems: since there are different types of stakeholders (e.g., data scientist, software engineer, and system user) of an ML system, they could make different but incompatible or invalid assumptions, leading to system misunderstanding, mismatch, etc lewis2021 . Zhang et al. Zhang2020 collected 4,960 real failures of several DL frameworks, such as TensorFlow, PyTorch, and Keras, and found that part of system failures are caused by assumptions. For example, during the evolution of TensorFlow and Keras, stakeholders assumed that Keras 2.2.2 can be used with TensorFlow 1.3 without being aware of the changes in the Softmax function in Keras 2.2.2, as shown in Figure 1, which led to system failures.
@@
Assumptions are related to many types of software artifacts, such as decisions, technical debt, and source code Yangsca2021 . For example, in TensorFlow, there is an assumption: “TODO: Looks like there is an assumption that weight has only one user. We should add a check here”, which induces a technical debt issue: a check should be added on whether the weight has a single user in the future. Existing research on assumptions and their management in software development usually uses experiments, surveys, and case studies to manually identify and extract assumptions through observation, questionnaires, interviews, focus groups, and documentation analysis Yangsms2018 . Since such approaches have high costs (e.g., time and resources), the comprehensiveness of the identified and extracted assumptions in these studies is often limited, leading to various problems in developing new theories, approaches, and methods of assumptions and their management in software development. In this work, to overcome the issues of manually identifying assumptions in DL framework development, we conducted an exploratory study to find a good algorithm and model to identify assumptions on GitHub. This provides a foundation for further research on assumptions and their management in the development of DL frameworks, for example, assumption classification, documentation, evaluation, and reasoning  Yangsms2018 .
@@
The remainder of the paper is organized as follows. Section 2 and Section 3 provide the related work and background knowledge, respectively, Section 4 presents the design of the evaluation on different classification models, Section 5 presents the results, Section 6 discusses the results, Section 7 clarifies the threats to the validity of the study, and Section 8 concludes the work with future directions.
@@ 2 Related work
In the field of assumptions and their management in software development, most assumptions are manually identified by researchers and practitioners. Landuyt and Joosen focused on the assumptions made during the application of a threat modeling framework (that is, LINDDUN), which allows the identification of privacy-related design flaws in the architecting phase Landuyt2020 . They conducted a descriptive study with 122 master students, and the students identified and extracted 845 assumptions from the models created by the students. In our previous work, we conducted an exploratory study as a first step in understanding assumptions in terms of their distribution, classification, and impacts based on code comments from nine popular DL framework projects on GitHub Yangsca2021 . The results show that (1) 3,084 assumptions are scattered across 1,775 files in the nine DL frameworks, ranging from 1,460 (TensorFlow) to 8 (Keras) assumptions. (2) There are four types of assumption validity: Valid, Invalid, Conditional, and Unknown and four types of assumptions based on their content: Configuration and Context, Design, Tensor and Variable, and Miscellaneous. (3) Both valid and invalid assumptions can have an impact within a specific scope (e.g., in a function) on the DL frameworks. Certain technical debt is induced when making assumptions, and source code is written and decisions are made based on assumptions. Xiong et al. studied the assumptions in the Hibernate Developer mailing list, including their expression, classification, trend over time, and related artifacts Xiong2018 . In their study, they identified and extracted 832 assumptions. Li et al. developed an ML approach Li2019 to identify and classify assumptions based on the dataset constructed by Xiong et al. Xiong2018 , which can read the data (i.e., sentences) from the dataset (i.e., a .csv file), preprocess the data (e.g., using NLTK and Word2Vec), train specific classifiers (e.g., Perception, Logistic Regression, and Support Vector Machine), and evaluate the trained classifiers (i.e., precision, recall, and F1 score).
@@
As described in the related studies mentioned above, assumptions and their management are important in the development of DL frameworks. However, existing approaches are either time-consuming or not specifically designed to identify assumptions in the development of DL frameworks.
@@ 3 Background
3.1 Machine learning algorithms
3.1.1 Perceptron
The Perceptron (Pct) algorithm is based on a threshold function as shown in Equation 1. $f(x)$ is the output, $w$ is a vector of weights, $x$ is the input, which is also a vector, and $b$ is the bias. The value of $f(x)$ is used to classify the input as type one or type two. Some studies used +1 and -1 instead of 1 and 0, which has the same rationale.
$f(x)=\begin{cases}1,ifw\cdot x+b>0\\ 0,otherwise\\ \end{cases}$ (1)
The Perceptron algorithm can also perform multiclass classification tasks, as shown in Equation 2. $x$ is the input, $y$ is the output, and $w$ is a vector of the weights. $f(x,y)$ is a feature representation function, which maps each possible input and output to a finite-dimensional feature vector. $\hat{y}$ is the resulting score, which represents the most possible of many outputs.
$\hat{y}=argmax_{y}f(x,y)\cdot w$ (2)
3.1.2 Logistic Regression
The Logistic Regression (LR) algorithm is similar to the Linear Regression algorithm, but uses a logistic function (that is, the sigmoid function) to deal with binary classification tasks, which is defined in Equation 3. $t$ is the input, and the output value is between 0 and 1 Jr2000 .
$\sigma(t)=\frac{1}{1+e^{-t}}$ (3)
Assuming that $t$ is linear, $x$ is the input vector, and $b$ is the bias (as shown in Equation 4), then the general logistic function can be described as Equation 5. $p(x)$ represents the probability that the output is in Category one and $(1-p(x))$ represents the probability that the output is in Category two.
$t=w\cdot x+b$ (4)
$p(x)=\sigma(t)=\frac{1}{1+e^{-(w\cdot x+b))}}$ (5)
LR can also be generalized to multiclass classification problems. Instead of using the Sigmoid function, LR can use the Softmax function to find the predicted probability of each category. For $i=1,....,N$, the standard Softmax function is shown in Equation 6. $N$ is the number of categories, $t$ is the input, and $t=(t_{1},t_{2},...,t_{N})\in\mathbb{R}^{N}$. The Softmax function uses an exponential function for each $t_{i}$ and ensures that the sum of the output vectors is 1.
$\sigma(t)_{i}=\frac{e^{t_{i}}}{\sum\limits_{j=1}^{N}e^{t_{j}}}$ (6)
3.1.3 Linear Discriminant Analysis
Linear Discriminant Analysis (LDA) can be used in both binary classification and multiclass classification Balakrishnama1998 . LDA maximizes the ratio of variance between classes to variance within classes in the datasets, which can ensure maximal separability Balakrishnama1998 . There are different strategies in LDA, for example: (1) maximize the ratio of variance between classes to variance within classes and (2) maximize the ratio of overall variance to variance within classes Balakrishnama1998 . Consider a binary classification problem with two datasets, the first step is to compute the mean (e.g., $\mu_{1}$ and $\mu_{2}$) of each dataset separately and the mean (e.g., $\mu_{3}$) of both datasets. Let $p_{1}$ and $p_{2}$ be the apriori probabilities of the classes, then we have Equation 7.
$\mu_{3}=p_{1}*\mu_{1}+p_{2}*\mu_{2}$ (7)
The within-class scatter is shown in Equation 8. For the binary problem, the within-class scatter is shown in Equation 9. Let $conv_{1}$ and $conv_{2}$ be the covariance of the training set and the test set. The equation of covariance is shown in Equation 10.
$S_{w}=\sum\limits_{j}p_{j}*(cov_{j})$ (8)
$S_{w}=p_{1}*cov_{1}+p_{2}*cov_{2}$ (9)
$cov_{j}=(x_{j}-\mu_{j})(x_{j}-\mu_{j})^{T}$ (10)
The between-class scatter is shown in Equation 11. Using different strategies in LDA will lead to a different calculation of the optimization criterion. After transformation of the data based on the selected strategy, we can further calculate the distance (e.g., the Euclidean distance) of the data from each class mean.
$S_{b}=\sum\limits_{j}(\mu_{j}-\mu_{3})*(\mu_{j}-\mu_{3})^{T}$ (11)
3.1.4 K-Nearest Neighbors
Fix and Hodges proposed a non-parametric method for pattern classification, which is the basis of K-Nearest Neighbors (KNN) Fix1989 . Cover and Hart expanded the KNN algorithm by adding formal properties of the rule Cover1967 . KNN can be used for both classification and regression problems. For the KNN classification, we first calculate the $k$ nearest neighbors of a data sample (e.g., the Euclidean distance). The data sample is then classified according to the vote of its $k$ nearest neighbors. Let $x_{i}=(x_{i}1,x_{i}2,…,x_{i}p)$ be a data sample. $p$ is the total number of features. The Euclidean distance between sample $x_{i}$ and $x_{j}(j=1,2,…,n)$ can be calculated using Equation 12.
$d(x_{i},x_{j})=\sqrt{(x_{i}1-x_{j}1)^{2}+(x_{i}2-x_{j}2)^{2}+...+(x_{i}p-x_{j}p)^{2}}$ (12)
3.1.5 Support Vector Machine
Support Vector Machine (SVM) is a commonly used supervised ML algorithm for data classification and regression analysis Cortes1995 . In a binary classification problem, SVM maps samples to a feature space and maximizes the width of the gap using a hyperplane between the two categories. If there is a new sample, it is mapped into the same feature space and assigned to Category 1 or 2 based on which side of the gap they fall into. SVM supports both linear classification and non-linear classification. For linear classification, taking into account $n$ points as $(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})$, $x_{i},i=1,2,...,n$ is a p-dimensional vector, $y_{i},i=1,2,...,n$ is the label of $x_{i}$, indicating which category $x_{i}$ belongs to. A hyperplane can be described as Equation 13. The parameter $\frac{b}{||w||}$ represents the offset of the hyperplane. For non-linear classification, SVM uses kernels to map the inputs into high-dimensional spaces, including Polynominal, Gaussian radial basis function, and Sigmoid function.
$w\cdot x-b=0$ (13)
3.1.6 Naive Bayes
Naive Bayes (NB) classifiers are based on the Bayes’ theorem and assume the features are independent McCallum1998 . Let the input $x$ have n features, which is $(x_{1},x_{2},...,x_{n})$. For Category $k$, the probability is $p(C_{k}|x)$. According to Bayes’ theorem, we can obtain Equation 14. NB classifiers usually use a decision rule with the model, i.e., the most probable one is the category.
$p(C_{k}|x)=\frac{p(C_{k})p(x|C_{k})}{p(x)}$ (14)
There are different types of NB models, such as the multivariate Bernoulli model and the multinomial model, which use different strategies to calculate the probability of each category McCallum1998 .
3.1.7 Classification and Regression Trees
In a classification problem, the non-leaf nodes of a decision tree are input features, and the leaf nodes of a decision tree are categories or probability distributions over the categories. The decision tree is constructed by iteratively splitting on the basis of specific rules. Considering there are three classes and two features, Figure 3 shows an example of the structure of a decision tree. A case goes to the left or right child node only if the condition is fulfilled. The leaf nodes represent the categories.
Figure 3: Structure of a decision tree Loh2011
There are many tree learning algorithms used for classification or regression problems, such as Iterative Dichotomiser 3 (ID3) Quinlan1986 , C4.5 (which is the successor of ID3) Quinlan1993 , Classification And Regression Trees (CART) Breiman1984 , and Chi-Square Automatic Interaction Detection (CHAID) Kass1980 . These algorithms use different strategies Loh2011 . For example, CART uses the Gini index to measure how well a node separates the data in order to choose the best split of the tree Loh2011 .
@@ 3.2 BERT and ALBERT
The architecture of the BERT model is based on the original Transformer model Vaswani2017 , which is a multilayer bidirectional Transformer encoder Devlin2019 , as shown in Figure 4. The BERT model was designed for multitasks, such as natural language understanding, question-and-answer, and sentence completion. In the original paper, Vaswani et al. proposed two versions of the BERT model: BERT-Base and BERT-Large Vaswani2017 . BERT-Base has 12 transformer blocks and 12 self-attention heads, the hidden size is 768, and the total parameters are 110M; BERT-Large has 24 transformer blocks and 16 self-attention heads, the hidden size is 1024, and the total parameters are 340M Vaswani2017 . To better deal with environments with restricted computational resources, Turc et al. proposed different versions of the BERT model with smaller sizes, for example, BERT-Tiny, BERT-Mini, BERT-Small, and BERT-Medium, which reduce the number of transformer blocks and hidden size turc2019 .
Figure 4: Architecture of the Transformer model Vaswani2017

### 20 Mar 2024 v3

@@ Abstract
Context: Stakeholders constantly make assumptions in the development of deep learning (DL) frameworks. These assumptions are related to various types of software artifacts (e.g., requirements, design decisions, and technical debt) and can turn out to be invalid, leading to system failures. Existing approaches and tools for assumption management usually depend on manual identification of assumptions. However, assumptions are scattered in various sources (e.g., code comments, commits, pull requests, and issues) of DL framework development, and manually identifying assumptions has high costs (e.g., time and resources).
Objective: The objective of the study is to evaluate different classification models for the purpose of identification with respect to assumptions from the point of view of developers and users in the context of DL framework projects (i.e., issues, pull requests, and commits) on GitHub.
Method: We constructed a new and largest dataset (i.e., AssuEval) of assumptions collected from the TensorFlow and Keras repositories on GitHub; explored the performance of seven traditional machine learning models (e.g., Support Vector Machine, Classification and Regression Trees), a popular DL model (i.e., ALBERT), and a large language model (i.e., ChatGPT) of identifying assumptions on the AssuEval dataset.
Results: The experiment results show that ALBERT achieves the best performance (f1-score: 0.9584) of identifying assumptions on the AssuEval dataset, which is much better than the other models (the 2nd best f1-score is 0.6211, achieved by ChatGPT). Though ChatGPT is the most popular large language model, we do not recommend using it to identify assumptions in DL framework development because of its low performance on the task. Fine-tuning ChatGPT specifically for assumption identification could improve the performance.
Conclusions: This study provides researchers with the largest dataset of assumptions for further research (e.g., assumption classification, evaluation, and reasoning) and helps practitioners better understand assumptions and how to manage them in their projects.
@@ 1 Introduction
The importance of assumptions and their management in software development (including DL framework development) has been highlighted in many studies. For example, Corbató Corbato1991 mentioned in his ACM Turing Award lecture that “design bugs are often subtle and occur by evolution with early assumptions being forgotten as new features or uses are added to systems.” Garlan et al. pointed out that incompatible assumptions in software architecture can cause architectural mismatch Garlan2009 . Lewis et al. also got similar results in ML systems: since there are different types of stakeholders (e.g., data scientist, software engineer, and system user) of an ML system, they could make different but incompatible or invalid assumptions, leading to system misunderstanding, mismatch, etc lewis2021 . Zhang et al. Zhang2020 collected 4,960 real failures of several DL frameworks, such as TensorFlow, PyTorch, and Keras, and found that part of system failures are caused by invalid assumptions. For example, during the evolution of TensorFlow and Keras, stakeholders assumed that Keras 2.2.2 can be used with TensorFlow 1.3 without being aware of the changes in the Softmax function in Keras 2.2.2, as shown in Figure 1, which led to system failures.
@@
Assumptions are related to many types of software artifacts, such as decisions, technical debt, and source code Yangsca2021 . For example, in TensorFlow, there is an assumption: “TODO: Looks like there is an assumption that weight has only one user. We should add a check here”, which induces a technical debt issue: a check should be added on whether the weight has a single user in the future. Existing research on assumptions and their management in software development usually uses experiments, surveys, and case studies to manually identify and extract assumptions through observation, questionnaires, interviews, focus groups, and documentation analysis Yangsms2018 . Since such approaches have high cost (e.g., time and resources), the comprehensiveness of the identified and extracted assumptions in these studies is often limited, leading to various problems in developing new theories, approaches, and methods of assumptions and their management in software development. In this work, to overcome the issues of manually identifying assumptions in DL framework development, we conducted an exploratory study to find a good algorithm and model to identify assumptions on GitHub. This provides a foundation for further research on assumptions and their management in the development of DL frameworks, for example, assumption classification, documentation, evaluation, and reasoning  Yangsms2018 .
@@
The remainder of the paper is organized as follows. Section 2 and Section 3 provide the related work and background knowledge, respectively, Section 4 presents the evaluation design for different classification models, Section 5 presents the results, Section 6 discusses the results, Section 7 clarifies the threats to the validity of the study, and Section 8 concludes the work with future directions.
@@ 2 Related work
In the field of assumptions and their management in software development, most assumptions are manually identified by researchers and practitioners. Landuyt and Joosen focused on the assumptions made during the application of a threat modeling framework (that is, LINDDUN), which allows the identification of privacy-related design flaws in the architecting phase Landuyt2020 . They conducted a descriptive study with 122 master students, and the students identified and extracted 845 assumptions from the models created by the students. In our previous work, we conducted an exploratory study as a first step to understand assumptions in terms of their distribution, classification, and impacts based on code comments from nine popular DL framework projects on GitHub Yangsca2021 . The results show that (1) 3,084 assumptions are scattered across 1,775 files in the nine DL frameworks, ranging from 1,460 (TensorFlow) to 8 (Keras) assumptions. (2) There are four types of assumption validity: Valid, Invalid, Conditional, and Unknown and four types of assumptions based on their content: Configuration and Context, Design, Tensor and Variable, and Miscellaneous. (3) Both valid and invalid assumptions can have an impact within a specific scope (e.g., in a function) on the DL frameworks. Certain technical debt is induced when making assumptions, and source code is written and decisions are made based on assumptions. Xiong et al. studied the assumptions in the Hibernate Developer mailing list, including their expression, classification, trend over time, and related artifacts Xiong2018 . In their study, they identified and extracted 832 assumptions. Li et al. developed an ML approach Li2019 to identify and classify assumptions based on the dataset constructed by Xiong et al. Xiong2018 , which can read the data (i.e., sentences) from the dataset (i.e., a .csv file), preprocess the data (e.g., using NLTK and Word2Vec), train specific classifiers (e.g., Perception, Logistic Regression, and Support Vector Machine), and evaluate the trained classifiers (i.e., precision, recall, and F1 score). We listed the differences of our work and the related work above in Table 1.
Table 1: Comparison of our work and the related work
Study Dataset Scale Identification Classification Models
Landuyt and Joosen Landuyt2020 Student projects 845 Manual Multiclass NA
Yang et al. Yangsca2021 DL frameworks 3,084 Manual Multiclass NA
Xiong et al. Xiong2018 Hibernate project 832 Manual Multiclass NA
Li et al. Li2019 Hibernate project 832 Automatic Binary ML models
Our work DL frameworks 15,354 Automatic Multiclass ML models, DL model, and LLM
@@
As described in the related studies mentioned above, assumptions and their management are important in the development of DL frameworks. However, existing approaches for assumption identification are typically either time-consuming or have been studied on a limited scale of data.
@@ 3.1 BERT and ALBERT
The architecture of the BERT model is based on the original Transformer model Vaswani2017 , which is a multilayer bidirectional Transformer encoder Devlin2019 . The BERT model was designed for multitasks, such as natural language understanding, question-and-answer, and sentence completion. In the original paper, Vaswani et al. proposed two versions of the BERT model: BERT-Base and BERT-Large Vaswani2017 . BERT-Base has 12 transformer blocks and 12 self-attention heads, the hidden size is 768, and the total parameters are 110M; BERT-Large has 24 transformer blocks and 16 self-attention heads, the hidden size is 1024, and the total parameters are 340M Vaswani2017 . To better deal with environments with restricted computational resources, Turc et al. proposed different versions of the BERT model with smaller sizes, for example, BERT-Tiny, BERT-Mini, BERT-Small, and BERT-Medium, which reduce the number of transformer blocks and hidden size turc2019 .
@@ 4 Study design
Specifically, we considered two types of assumptions: self-claimed assumptions (SCAs) and potential assumptions (PAs). The rationale is that a sentence cannot be identified as an assumption unless it is explicitly mentioned by stakeholders, which are SCAs. There are also some plausible cases that may or may not be assumptions, which we have to ask stakeholders to confirm, which are PAs. Taking into account a sentence in a commit, PR, or issue of a GitHub repository, an SCA is that the sentence includes an assumption and the assumption is explicitly claimed using at least one of the assumption-related terms (i.e., “assumption”, “assumptions”, “assume”, “assumes”, “assumed”, “assuming”, “assumable”, and “assumably”). For example, in this sentence “[tf/xla] fixup numbering of xla parameters used for aliasing previously, the xla argument parameter was incorrectly assumed to correspond to the index in the vector of ‘xlacompiler::argument’”, it includes an SCA of assuming that the xla argument parameter corresponds to the index in the vector of xlacompiler::argument and the developer claimed that it was an invalid assumption. A PA is that the sentence may include an assumption, which is a potential assumption that needs further confirmation from human experts. This definition covers various aspects, such as expectations, future events, possibilities, guesses, opinions, feelings, and suspicions, which can indicate PAs. We provide three examples to further explain the concept of PA. For example, in the sentence “I think the right way to create demo tensorboard instances is to simply run a tensorboard in the cloud, rather than keep maintaining this mocked-out backend.”, it includes a PA regarding thoughts on the right way to create demo tensorboard instances. In another example “The system will not crash under heavy load”, the sentence describes the future state of the system, which is uncertain and includes a PA. The third example “both false and true outputs should be considered independently” does not include assumption-related terms, but the sentence describes an expectation (i.e., “something should be”), which is a PA. After further confirmation by human experts, this PA can be transformed into an SCA or other types of software artifacts. Besides SCAs and PAs (which belong to explicit assumptions), there are also many implicit assumptions in projects (e.g., in stakeholders’ heads or requiring reasoning). Identifying implicit assumptions is much more tricky than identifying SCAs and PAs, since there are no explicit clues in the sentences and stakeholders need to infer the sentences based on the context, which involves assumption reasoning. Identifying implicit assumptions is beyond the scope of this study and we treat it as future work. Therefore, for a sentence that is neither an SCA nor a PA, we consider it as “not an assumption” (NA).
In this study, we used and analyzed seven traditional ML models (e.g., Support Vector Machine, Classification and Regression Trees), a popular DL model (i.e., ALBERT), and an LLM (i.e., ChatGPT) to conduct the study. For some cases, BERT and ALBERT may also be considered as LLMs. We did not train the ALBERT model with a large corpus of text, and therefore we treated the ALBERT model as a DL model instead of an LLM. Additionally, rule-based approaches may be suitable for identifying SCAs (i.e., searching radical “assum” in sentences), but we see the following flaws: (1) There are many variables, functions, etc. named with the radical “assum” (e.g., “assume”), which is difficult to filter out in all cases. (2) Sentences may have terms with the radical “assum”, while the sentences are about whether to make certain assumptions (e.g., “@willnorris for corporate CLAs, do we expect googlebot to say this has been signed? or do we just assume that if the user has @dropbox.com in their email that it is fine?”), that is, the assumptions do not exist. Using a simple rule-based approach may misclassify such cases as SCAs, which are NAs. (3) Rule-based approaches may not be suitable for the identification of PAs based on the definition. For example, consider the following two sentences that have the term “possible”: a. “Along with the template, please provide as many details as possible to find the root cause of the issue like CUDA and cuDNN versions.” and b. “@YANxu-666: we don’t have any continuous integration for Windows so it is possible that something in the script is not Windows-compatible.”, Sentence a is clearly not an assumption while Sentence b is a PA. In such a case, considering phrases instead of words may address the problem. However, there are many words that can lead to assumptions, and the phrases can add much more complexity to the assumption identification algorithm. Therefore, in this study, we did not analyze the performance of rule-based approaches in assumption identification.